Let X be continuous with pdf(x) = 3x^2 if 0 < x < 1, and zero otherwise,

a) Find E(X).

b)Find Var(X).

c)Find E(X').

d)Find E(3X - 5X^2 + 1).


1
Expert's answer
2021-11-09T16:29:22-0500

a)


"E(X)=\displaystyle\int_{-\infin}^{\infin}xf(x)dx=\displaystyle\int_{0}^{1}x(3x^2)dx"

"=\dfrac{3}{4}"

b)


"E(X^2)=\displaystyle\int_{-\infin}^{\infin}x^2f(x)dx=\displaystyle\int_{0}^{1}x^2(3x^2)dx"

"=\dfrac{3}{5}"

"Var(X)=E(X^2)-(E(X))^2=\dfrac{3}{5}-(\dfrac{3}{4})^2"

"=\dfrac{3}{80}"

c)


"E(X^r)=\displaystyle\int_{-\infin}^{\infin}x^rf(x)dx=\displaystyle\int_{0}^{1}x^r(3x^2)dx"

"=\dfrac{3}{r+3}, r\not=-3"


"E(X^{-3})=\displaystyle\int_{-\infin}^{\infin}x^{-3}f(x)dx=\displaystyle\int_{0}^{1}x^{-3}(3x^2)dx"

"=[3\ln(|x|)]\begin{matrix}
 1 \\
 0
\end{matrix}=does \ not\ exist"

d)


"E(3X-5X^2+1)=3E(X)-5E(X^2)+1"

"=3(\dfrac{3}{4})-5(\dfrac{3}{5})+1=\dfrac{1}{4}"